Which of the following oxoacids has the smallest value for Ka,1 ?

H2TeO3

H2SO3

H2TeO4

H2SeO4

H2SeO3

H2SO4

Which of the following choices is appropriate if you have made an HBrO/NaBrO buffer and the pH is 8.0, and you want to have the buffer with equal amounts of conjugate components. Choose all that apply.

Add some strong base

Add some strong acid

Add some weak base

Add some weak acid

Add some water
Answer #1

Answer! o Answer: The and having smallest value of Kai H₂ Te Oz Explanations : la fello-h HO 611 This ease of release of ProtHoBoo/NaBro buffer , pn=8-0 Pka af HB00= 8.64 PM af this acides buffer is given by pre pka + log na Breg 8.0 = 8.64 + log [Na